OpenStudy (anonymous):

evaluate the limit as x goes to 0: ((x(3+x))/(6x^2)

OpenStudy (anonymous):

1/6

OpenStudy (amistre64):

the trick here is to see if we can cancel the part that is giving troubles

OpenStudy (anonymous):

u can use l'hospital though :)

OpenStudy (amistre64):

\[\frac{x(3+x)}{6.x.x}\]

OpenStudy (anonymous):

aaaagh! that is what I missed! Thank you!

OpenStudy (amistre64):

\[\frac{3+x}{6x}\] \[\frac{3}{6x}+\frac{x}{6x}\]

OpenStudy (anonymous):

that was a tricky one - thanks for explaining so well!!!

OpenStudy (anonymous):

i/1 though is not one of my choices?

OpenStudy (anonymous):

1/6

OpenStudy (anonymous):

you will always have an undefined in the denom anyway

OpenStudy (amistre64):

what are your choices?

OpenStudy (anonymous):

1/2 DNE -3 0 3

OpenStudy (anonymous):

must be DNE

OpenStudy (amistre64):

its DNE; the one sided limits dont agree

OpenStudy (anonymous):

that's what I got yay!

OpenStudy (amistre64):

http://www.wolframalpha.com/input/?i=lim {x+to+0}%28%28x%283%2Bx%29%29%2F%286x^2%29 for clarification :)

OpenStudy (amistre64):

it cuts the link short .... bummer

OpenStudy (anonymous):

thanks anyway

OpenStudy (anonymous):

nope for this u should not split and evaluate the limit... u need to apply the l'hospital rule directly, this'll giv u the correct answer.